a) Find the distance from a point charge q=100nC where the field intensity is equal to E=6kN/C

b) The charge density on each of two parallel plates is of 5μc/m2. Find the intensity of the electric field E between both plates.



1
Expert's answer
2021-10-01T13:27:19-0400

a) Write the expression for field and determine the distance:


"E=\frac{kq}{r^2},\\\space\\
r=\sqrt{kq/E}=0.39\text{ m}."

b) The field is


"E=\frac\sigma{2\epsilon_0}=282\text{ kN/C}."